The perimeter of a rectangle is 34ft, the length is 3ft longer than the width. Find the dimension

Found 2 solutions by zeti68, richwmiller:
Answer by zeti68(10)

Hi,
length=L
width=W
34= 2(L+W)
L= W+3
34= 2(W+3+W)= 2(2W+3)= 4W+6
34-6=4W
28= 4W
28/4=W
W= 7
L= 7+3
L=10
Answer: length=7 ft
width=10 ft

Answer by richwmiller(17219)

34-6=28
width is 7
length is 10
check
2*10+2+7=34
traditional way
l=w+3
2w+2l=34
w+l=17
w+w+3=17
2w=14
w=7
l=10